# On endosplit $p$-permutation resolutions and Broué’s conjecture for $p$-solvable groups

### v2

@@ 5. Constructing splendid Rickard equivalences for blocks of $p$-solvable groups over $\mathbb{F}_{p}$
Let $G$ be a $p$-solvable group satisfying $G=O_{p^{\prime},p,p^{\prime}}(G)$ with a Sylow $p$-subgroup $P$. Let $b$ be a $G$-stable block idempotent of $\mathbb{F}_{p}O_{p^{\prime}}(G)$. Then $b$ is a block of $\mathbb{F}_{p}G$, and if $c$ is its Brauer correspondent, then $b$ and $c$ are splendidly Rickard equivalent.
@@
Proposition 5.8 implies $b$ is a block of $\mathbb{F}_{p}G$. Let $b^{\prime}$ be a block of $\overline{\mathbb{F}}_{p}O_{p^{\prime}}(G)$ for which $bb^{\prime}\neq 0$, then from Proposition 5.8 $b^{\prime}$ is a block of $G$ which is again $G$-stable, as $b^{\prime}$ is a trace sum of Galois conjugates of $b$. Let $c^{\prime}$ be the Brauer correspondent of $b^{\prime}$, in which case we have $cc^{\prime}\neq 0$. Then [HL00] proves $b^{\prime}$ and $c^{\prime}$ are splendidly Rickard equivalent, and Remark 5.9 demonstrates that the equivalence may be realized over some finite field $k$. Construction 5.11 verifies that the construction of [HL00] descends to $k[b]=k[c]$, and [KL18, Theorem 6.5(b)] implies that $b^{\prime}$ and $c^{\prime}$ are splendidly Rickard equivalent over $\mathbb{F}_{p}$ (after using Rickard’s lifting theorem [Ric96, Theorem 5.2] to lift the equivalence of $b$ and $c$ to an absolutely unramified complete discrete valuation ring). ∎
As an immediate corollary, we obtain another proof that various refinements of the Alperin-McKay conjecture proposed by Issacs-Navarro in [IN02], Navarro in [Nav04], and Turull in [Tur13] hold in our setting. We refer the reader to [Bol24, Page 5] for an overview of these refinements of the Alperin-McKay conjecture. Note that Turull’s reformulation, the strongest of the three refinements, was shown by Turull in [Tur13] for all $p$-solvable groups using separate methods.
@@
Refinements (5), (7), and (8) of the Alperin-McKay conjecture, as given in [Bol24], hold for blocks of finite $p$-solvable groups $G$ satisfying $G=O_{p^{\prime},p,p^{\prime}}(G)$ with abelian defect group.
@@
This is an immediate corollary of [Bol24, Theorem 1.4], after applying unique lifting of splendid Rickard equivalences from $k$ to $\mathcal{O}$, since splendid Rickard equivalences induce $p$-permutation equivalences (see [BX08, Theorem 1.5]). ∎
@@
The key obstruction to verifying Kessar and Linckelmann’s conjecture for all $p$-solvable groups using the techniques here is that [HL00, Lemma 3.3] may not hold for fields not large enough for $G$. Indeed, a counterexample is as follows. Consider $\mathbb{F}_{2}S_{3}=\mathbb{F}_{2}(C_{3}\rtimes C_{2})$, which has two blocks. The non-principal block idempotent $e$ is a block of $\mathbb{F}_{2}C_{3}$. However, in $\mathbb{F}_{4}C_{3}$, $e$ decomposes as $e=e_{1}+e_{2}$, where the two blocks $e_{1},e_{2}$ correspond to the two nontrivial characters of $C_{3}$, and $C_{2}$ swaps $e_{1}$ and $e_{2}$. It follows that $e$ is a $S_{3}$-stable block of $\mathbb{F}_{2}O_{2^{\prime}}(S_{3})=\mathbb{F}_{2}C_{3}$, but $e$ has trivial defect group, not defect group the Sylow 2-subgroup $C_{2}.$ The author thanks the anonymous referee for an earlier version of this paper for providing this counterexample.
Acknowledgements: The author thanks Markus Linckelmann for his assistance which helped the author understand certain arguments in [HL00], and the anonymous referee for pointing out a faulty lemma and providing an explicit counterexample for a previous version of this paper.

### v3

@@ 5. Constructing splendid Rickard equivalences for blocks of $p$-solvable groups over $\mathbb{F}_{p}$
Let $G$ be a $p$-solvable group satisfying $G=O_{p^{\prime},p,p^{\prime}}(G)$ with a abelian Sylow $p$-subgroup $P$. Let $b^{\prime}$ be a $G$-stable block of $\overline{\mathbb{F}}_{p}O_{p^{\prime}}(G)$, and let $b$ be a block idempotent of $\mathbb{F}_{p}G$ for which $bb^{\prime}\neq 0$. Let $c$ be the Brauer correspondent of $b$ in $\mathbb{F}_{p}N_{G}(P)$. Then $b$ and $c$ are splendidly Rickard equivalent.
@@
Proposition 5.8 implies $b^{\prime}$ is a block of $\overline{\mathbb{F}}_{p}G$ with defect group $P$. Let $c^{\prime}$ be the Brauer correspondent of $b^{\prime}$, in which case we have $cc^{\prime}\neq 0$ as well. Then [HL00] proves $b^{\prime}$ and $c^{\prime}$ are splendidly Rickard equivalent, and Remark 5.9 demonstrates that the equivalence may be realized over some finite field $k$. Construction 5.11 verifies that the construction of [HL00] descends to $k[b]=k[c]$, and [KL18, Theorem 6.5(b)] implies that $b$ and $c$ are splendidly Rickard equivalent over $\mathbb{F}_{p}$ (after using Rickard’s lifting theorem [R96, Theorem 5.2] to lift the equivalence of $b$ and $c$ to an absolutely unramified complete discrete valuation ring). ∎
If $G=O_{p^{\prime},p,p^{\prime}}(G)$ with abelian Sylow $p$-subgroup, we say a block idempotent $b$ of $\mathbb{F}_{p}G$ satisfies condition $(\ast)$ if it arises in the situation of Theorem 5.12. That is, there exists a $G$-stable block $b^{\prime}$ of $\overline{\mathbb{F}}_{p}O_{p^{\prime}}(G)$ satisfying $bb^{\prime}\neq 0$. As an immediate corollary, we obtain another proof that various refinements of the Alperin–McKay conjecture proposed by Isaacs–Navarro in [IN02], Navarro in [N04], and Turull in [T13] hold in our setting. We refer the reader to [Bo22, Page 5] for an overview of these refinements of the Alperin–McKay conjecture. Note that Turull’s reformulation, the strongest of the three refinements, was shown by Turull in [T13] for all $p$-solvable groups using separate methods.
@@
Refinements (5), (7), and (8) of the Alperin–McKay conjecture, as given in [Bo22], hold for blocks satisfying condition $(\ast)$.
@@
This is an immediate corollary of [Bo22, Theorem 1.4], after applying unique lifting of splendid Rickard equivalences from $k$ to $\mathcal{O}$, since splendid Rickard equivalences induce $p$-permutation equivalences (see [BX07, Theorem 1.5]). ∎
@@
The key obstruction to verifying Kessar–Linckelmann’s conjecture for all $p$-solvable groups by proving the reduction step [HL00, Proposition 3.1] is that [HL00, Lemma 3.3] may not hold for fields not large enough for $G$. Indeed, a counterexample is as follows. Consider $\mathbb{F}_{2}S_{3}=\mathbb{F}_{2}(C_{3}\rtimes C_{2})$, which has two blocks. The non-principal block idempotent $e$ is a block of $\mathbb{F}_{2}C_{3}$. However, in $\mathbb{F}_{4}C_{3}$, $e$ decomposes as $e=e_{1}+e_{2}$, where the two blocks $e_{1},e_{2}$ correspond to the two nontrivial characters of $C_{3}$, and $C_{2}$ swaps $e_{1}$ and $e_{2}$. It follows that $e$ is a $S_{3}$-stable block of $\mathbb{F}_{2}O_{2^{\prime}}(S_{3})=\mathbb{F}_{2}C_{3}$, but $e$ has trivial defect group, not defect group the Sylow 2-subgroup $C_{2}.$
This also demonstrates that not all block idempotents of $\mathbb{F}_{p}G$, where $G$ satisfies $G=O_{p^{\prime},p,p^{\prime}}(G)$ with abelian Sylow $p$-subgroup, satisfy condition $(\ast)$. Indeed, with the above notation, there does not exist a $G$-stable block $e^{\prime}$ of $\overline{\mathbb{F}}_{2}C_{3}$ such that $ee^{\prime}\neq 0$, as $e_{1}$ and $e_{2}$ are not $G$-stable.
Acknowledgments: The author thanks Markus Linckelmann for his assistance which helped the author understand certain arguments in [HL00], and a referee for pointing out a faulty lemma and providing the above counterexample for a previous version of this paper, which had attempted to prove [HL00, Proposition 3.1] over an arbitrary field of characteristic $p$.